40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 322 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  1792   Sat Jul 25 19:04:01 2009 KojiUpdatePSLAligning the beam to the Faraday

Quote:

Quote:

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

 

At least one problem is the mis-centering of the resonant spot on MC2, which can be viewed with the video monitors.  It's very far from the center of the optic, which causes length-to-angle coupling that makes the mulitple servos which actuate on MC2 (MCL, WFS, local damping) fight each other and go unstable.

I played with the MC alignment for the beam centering. After that, I restored the alignment values.



In principle, one can select the MC2 spot as one likes, while the transmitted beam axis to the IFO is not changed
as far as you are at the best alignment. This principle is almost trivial because the beam axis matches
to the input beam axis at the best alignment.
The alignment solution is not unique for a triangle cavity if we don't fix the end spot position.

In practice, this cruising of the MC2 spot is accomplished by the following procedure:
0) Assume that you are initially at the best alignment (=max transmission).
1) Slightly tilt the MC2.
2) Adjust MC1/MC3 so that the best transmission is restored.

I started from the following initial state of the alignment sliders:

BEFORE TRIAL

MC1 Pitch  +3.6242
MC1 Yaw  -0.8640
MC2 Pitch  3.6565
MC2 Yaw -1.1216
MC3 Pitch -0.6188
MC3 Yaw -3.1910
MC Trans 2.70

After many iterations, the spot was centered in some extent. (See the picture)
RESULT

    adj.
MC1 Pitch  +3.363 (-0.26)
MC1 Yaw  -1.164 (-0.3)
MC2 Pitch  3.7565 (+0.1)
MC2 Yaw -1.2800 (~ -0.16)
MC3 Pitch -0.841 (~ -0.22)
MC3 Yaw -3.482 (~ -0.29)
MC Trans 2.75  

The instability looked cured somewhat.
Further adjustment caused a high freq (10Hz at the camera) instability and the IMCR shift issue.
So I returned to the last stable setting.

Side effect:
Of course, if you move MC1, the reflected spot got shifted.
The spot has been apparently off-centered from the IMCR camera. (up and right)
At this stage, I could not determine what is the good state.
So, I restored the alignment of the MC as it was.
But now Alberto can see which mirror do we have to move in which direction and how much.

Attachment 1: MC2_Cam.jpg
MC2_Cam.jpg
  1791   Sat Jul 25 16:04:32 2009 robUpdatePSLAligning the beam to the Faraday

Quote:

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

 

At least one problem is the mis-centering of the resonant spot on MC2, which can be viewed with the video monitors.  It's very far from the center of the optic, which causes length-to-angle coupling that makes the mulitple servos which actuate on MC2 (MCL, WFS, local damping) fight each other and go unstable.

  1790   Sat Jul 25 13:49:28 2009 KojiUpdateGeneralMultiply Resonant EOM Update

Quote:

After speaking with Rana and realizing that it would be better to use smaller inductances in the flying-component circuit (and after a lot of tinkering with the original), I rebuilt the circuit, removing all of the resistors (to simplify it) and making the necessary inductance and capacitance changes. A picture of the circuit is attached, as is a circuit diagram.

A plot of the measured and simulated transfer functions is also attached; the general shape matches between the two, and the resonant frequencies are roughly correct (they should be 11, 29.5, and 55 MHz). The gain at the 55 MHz peak is lower than the other two peaks (I'd like them all to be roughly the same). I currently have no idea what the impedance is doing, but I'm certain it is not 50 Ohms at the resonant peaks, because there are no resistors in the circuit to correct the impedance. Next, I'll have to add the resistors and see what happens.

Stephanie, 

This is a quite nice measurement. Much better than the previous one.

1) For further steps, I think now you need to connect the real EOM at the end in order to incorporate
the capacitance and the loss (=resistance) of the EOM. Then you have to measure the input impedance
of the circuit. You can measure the impedance of the device at Wilson house.
(I can come with you in order to consult with Rich, if you like)

Before that you may be able to do a preparatory measurement which can be less precise than the Wilson one,
but still useful. You can measure the transfer function of the voltage across the input of this circuit,
and can convert it to the impedance. The calibration will be needed by connecting a 50Ohm resister
on the network analyzer.

2) I wonder why the model transfer function (TF) has slow phase changes at the resonance.
Is there any implicit resistances took into account in the model?

If the circuit model is formed only by reactive devices (Cs and Ls), the whole circuit has no place to dissipate (= no loss).
This means that the impedance goes infinity and zero, at the resonance and the anti-resonance, respectively.
This leads a sharp flip of the phase at these resonances and anti-resonances.

The real circuit has small losses everywhere. So, the slow phase change is reasonable.

  1789   Sat Jul 25 13:34:58 2009 KojiUpdateGeneralWeek 5/6 Update

Quote:

The last week I've started setting up the HeNe laser on the PSL table and doing some basic measurements (Beam waist, etc) with the beam scan, shown on the graph.  Today I moved a few steering mirrors that steve showed me from at table on the NW corner to the PSL table.  The goal setup is shown below, based on the UCSD setup.  Also, I found something that confused me in the EUCLID setup, a  pair of quarter wave plates in the arm of their interferometer, so I've been working out how they organized that to get the results that they did.  I also finished calculating the shot noise levels in the basic and UCSD models, and those are also shown below (at 633nm, 4mw) where the two phase-shifted elements (green/red) are the UCSD outputs, in quadrature (the legend is difficult to read).

 

 

Chris,

Some comments:

0. Probably, you are working on the SP table, not on the PSL table.

1. The profile measurement looks very nice.

2. You can simplify the optical layout if you consider the following issues
  A. The matching lenses just after the laser:
      You can make a collimated beam only with a single lens, instead of two.
      Just put a lens of f0 with distance of f0 from the waist. (Just like Geometrical Optics to make a parallel-going beam.)

      Or even you don't need any lens. In this case, whole optical setup should be smaller so that your beam
      can be accomodated by the aperture of your optics. But that's adequately possible.

  B. The steering mirrors after the laser:
      If you have a well elevated beam from the table (3~4 inches), you can omit two steering mirrors.
      If you have a laser beam whose tilte can not be corrected by the laser mount, you can add a mirror to fix it.

  C. The steering mirrors in the arms:
      You don't need the steering mirrors in the arms as all d.o.f. of the Michelson alignment can be adjusted
      by the beamsplitter and the mirror at the reflected arm. Also The arm can be much shorter (5~6 inches?)

  D. The lenses and the mirrors after the PBS:
      You can put one of the lenses before the PBS, instead of two after the lens.
      You can omit the mirror at the reflection side of the PBS as the PBS mount should have alignment adjustment.

The simpler, the faster and the easier to work with!
Cheers.

  1788   Fri Jul 24 21:02:46 2009 AlbertoUpdatePSLAligning the beam to the Faraday

This afternoon I kept working on the alignment of the beam so that it matches at the same the PSL periscope, the Mode Cleaner and the Faraday isolator at the input of the IFO.

The camera looking at the Farady showed a beam quite low from the center of the Faraday's entrance. I wanted to move it up.

After working on the periscope alignment and on the MC mirrors, I think I managed to moved it up a bit. To know whether that was enough or not I wanted to evaluate the alignment to the X arm by checking the value of TRX.

In order for the MC to be finely matched to the input beam from the periscope, the WFS controls have to be on. Before turning them on, I centered the beam on their QPDs and run the WFS_zero_offset script.

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

Flashes at ETMX show at least that the beam is going through the Farady. How well, I can't tell untill the MC is under full control.

 

I have to leave the lab now, but I can be back tomorrow to keep working on that.

  1787   Fri Jul 24 17:47:52 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

After speaking with Rana and realizing that it would be better to use smaller inductances in the flying-component circuit (and after a lot of tinkering with the original), I rebuilt the circuit, removing all of the resistors (to simplify it) and making the necessary inductance and capacitance changes. A picture of the circuit is attached, as is a circuit diagram.

A plot of the measured and simulated transfer functions is also attached; the general shape matches between the two, and the resonant frequencies are roughly correct (they should be 11, 29.5, and 55 MHz). The gain at the 55 MHz peak is lower than the other two peaks (I'd like them all to be roughly the same). I currently have no idea what the impedance is doing, but I'm certain it is not 50 Ohms at the resonant peaks, because there are no resistors in the circuit to correct the impedance. Next, I'll have to add the resistors and see what happens.

Attachment 1: BuiltCkt2_Picture_Simplified.png
BuiltCkt2_Picture_Simplified.png
Attachment 2: BuiltCkt2_Simplified.png
BuiltCkt2_Simplified.png
Attachment 3: Simplified.png
Simplified.png
  1786   Fri Jul 24 17:20:48 2009 JenneUpdateoplevsETMY oplev is iffy

ETMY oplev is currently a work in progress.  The HeNe beam is hitting the photodiode, but the spot size there is pretty much the size of the entire QPD.  Thus, the ETMY oplev isn't really useful right now.  I'm re-figuring things out (note to self: close to the laser, you have to use Gaussian optics...regular ray tracing doesn't really work), and hopefully will have the oplev back under control by the time Alberto is finished realigning the IFO, so this doesn't keep anyone from doing any exciting locking work.

  1785   Fri Jul 24 11:04:11 2009 AlbertoConfigurationComputerselog restarted

This morning I found the elog down. I restarted it using the procedure in the wiki.

  1784   Thu Jul 23 20:30:23 2009 AlbertoUpdatePSLBeam aligned to the Farady

After yesterday's changes in the MC cavity state today it was necessary to optimize the alignment to the Faraday.

The way I did it was by tuning the PSL periscope in pitch and yaw trying to maximize TRX with the arm locked. After a small change in either one of the two directions I first maximized the MC transmitted power and then I ran the alignment script for the X arm.

I explored the space for both pitch and yaw and the max that I could get from TRX was 0.91. I'm not sure whether the increase in TRX is entirely due to a better alignment to the Farady rather than to a higher MC transmitted power.

Also I'm not sure I'm well interpreting the image from the camera pointing at the Farady. I guess I need someone more familiar with it to tell me if it shows any sign of clipping.

Anyway, last week, even before the MC got misaligned, TRX didn't go above 0.90. So now I wonder whether it's the MC's fault or something else's if we have that value..

  1783   Thu Jul 23 10:05:38 2009 AlbertoUpdatePSLSummary of the latest adventures with the alignment of the mode cleaner

Alberto, Koji,

Summary of the latest adventures with the alignment of the mode cleaner

Prior events.

  • Last week, on July 12th the RFM network crashed (elog entry 1736). I don't know for sure which one was the cause and which one the effect, but also C1DAQADW was down and it didn't want to restart. Alex fixed it the day after.
  • On the evening of Sunday July 20th I noticed that the mode cleaner was unlocked. A closer inspection showed me that MCL was frozen at -32768 counts. To fix that I rebooted C1DCUEPICS and burtrestored to snapshots from the day before.
  • On Tuesday July 21st another failure of the RFM Network made necessary a reboot of the frame builder and of all front end computers (entry 1772). As a consequence, the mode cleaner couldn't get locked anymore, even if the mirror's sliders in the MC-Align MEDM screen were in the proper positions. At that time I missed to check the MC suspension positions as a way to ensure that the MC hadn't really changed. Although later, as it turned out, that would have been useless anyway since all the data record prior to the computers crash of that day somehow had been corrupted (entry 1774). Neither the MC2 LSC control or MC ASC control could engage so I (erroneously) thought that some tune of the periscope might help. So I did but, since the Mode Cleaner was misaligned, that had the effect of spoiling the good matching of the periscope to the MC cavity.
  • Yesterday, Wednesday July 22nd, I found out about the sticky slider effect (entry 1776). At that point we didn't have anymore a way to know that the MC optics were actually in their proper original alignment state because of the lack of a reference for those in the data record (as I wrote above). I had to go back to the periscope and fix the alignment.


Chronicles of periscope and MC alignment

Yesterday morning I started aligning the periscope but it turned out to be trickier than usual. With the ASC (Alignment Sensing Control) off and only the length controls on, the Mode Cleaner didn't lock easily, although I knew I wasn't very far from the sweet spot.

In the afternoon the struggle continued and the matching of the the beam to the MC cavity became just worse. At some point I noticed that the ASC inputs somehow had got on - although the ASC still looked disabled from the MClock MEDM main screen. So I was actually working against the Wave Front Sensors and further worsening the periscope alignment.

That hurled me to the weeds. After hours of rowing across the stormy waters of a four-dimensional universe I got to have occasional TEM00 flashes at the transmission but still, surprisingly, no MC locking. Confused, I kept tuning the periscope but that just kicked me off road again.

Then at about 7pm Koji came to my rescue and suggested a more clever and systematic way to solve the problems. He suggested to keep record of the MC mirrors alignment state and re-align the cavity to the periscope. Then we would gradually bring the cavity back to the original good position changing the periscope alignment
at the same time.

 

That would have worked straight away, if we hadn't been fighting against a subtle and cruel enemy: the 40m computer network. But I (as John Connor), and Koji (as the Terminator) didn't pull back.

Here's a short list of the kinds of weapons that the computers threw to us:

  1. After a while the FSS entered a funny state. It showed transmission: we had light at the MC (and even flashes) but the MEDM readout of the FSS transmitted power after the cavity was low (~0.019). Also the spot on the monitor showed a slightly different pattern from how I remembered it. On the other side the transmission camera didn't show that typical halo as usual.
  2. MCL was frozen at 32768. I ran the MCDown and MCUp script a couple of times and that unstuck it.
  3. On op340m we found that the MC autolocker script wasn't running. So I restarted it. Still nothing changed: bright and sharp flashes appeared on the monitor (sign of a not too bad alignment) but no lock.
  4. I rebooted C1IOO. No change.
  5. I rebooted C1DCUEPICS and burtrestored the EPICS computers to Jul 19th. No change.
  6. Then I burtrestored the c1psl.snapshot and that finally did something. The FSS reflected spot changed and the halo appeared again at its transmission camera. Soon after the MC got locked.


We then proceeded with Koji's plan. In an iterative process, we aligned the MC cavity maximizing the transmission and tuned the periscope in order to match the Faraday input of the interferometer. The last thing we did it by looking at the camera pointing at the Faraday isolator.

We found that we didn't have to tune the periscope much. That means that all afternoon I didn't really go too far, but the autolocker wasn't working properly, or it wasn't working at all.

Then we ran the alignment script for the X arm but it didn't work before we aligned the steering mirrors.

Then we ran it three times but could not get more than 0.87 at TRX. That means that there we still have to work on the alignment to the Faraday. That's job for today in the trenches of the lab.

 

  1782   Thu Jul 23 07:34:45 2009 AidanUpdateCDSAdded C2 MEDM screens to 40m SVN.

 

See Adhikari eLOG entry: http://nodus.ligo.caltech.edu:8080/AdhikariLab/194

  1781   Wed Jul 22 20:11:26 2009 peteUpdateComputersRCG front end

I compiled and ran a simple (i.e. empty) front end controller on scipe12 at wilson house.  I hooked a signal into the ADC and watched it in the auto-generated medm screens. 

There were a couple of gotchas:

1. Add an entry SYS to the file /etc/rc.local, to the /etc/setup_shmem.rtl line, where the system file is SYS.mdl.

2. If necessary, do a BURT restore.  Or in the case of a mockup set the BURT Restore bit (in SYS_GDS_TP.adl) to 1.

 

  1780   Wed Jul 22 18:04:14 2009 robOmnistructureComputersweird noise coming from Gigabit switch

in the rack next to the printer.  It sounds like a fan is hitting something.

  1779   Wed Jul 22 16:15:52 2009 Chris ZimmermanUpdateGeneralWeek 5/6 Update

The last week I've started setting up the HeNe laser on the PSL table and doing some basic measurements (Beam waist, etc) with the beam scan, shown on the graph.  Today I moved a few steering mirrors that steve showed me from at table on the NW corner to the PSL table.  The goal setup is shown below, based on the UCSD setup.  Also, I found something that confused me in the EUCLID setup, a  pair of quarter wave plates in the arm of their interferometer, so I've been working out how they organized that to get the results that they did.  I also finished calculating the shot noise levels in the basic and UCSD models, and those are also shown below (at 633nm, 4mw) where the two phase-shifted elements (green/red) are the UCSD outputs, in quadrature (the legend is difficult to read).

 

 

Attachment 1: Beam_Scan.jpg
Beam_Scan.jpg
Attachment 2: Long_Range_Michelson_Setup_1_-_Actual.png
Long_Range_Michelson_Setup_1_-_Actual.png
Attachment 3: NSD_Displacement.png
NSD_Displacement.png
  1778   Wed Jul 22 14:44:57 2009 ZachUpdateCamerasGigE Phase Camera

This past week, I have mostly been debugging my software.  I have tried to use the fluorescent lights to test the camera, but I can't tell for sure if my code is finding the correct amplitude and phase or not.  I am currently using Mathematica to double check my calculations in solving for the phase and amplitude.

Also, I have taken dark field images using a lens with a closed shutter.  I have found that the dark band across the top of the images only appears after the camera heats up.  Also, there is an average electronic noise of 14 with a maximum of 40.  However, this electronic noise as well as any consistent ambient noise will be automatically corrected for in the calculations I'm using because I'm taking the differences between the CCD images to calculate relative phases and amplitudes.

I should be able to start setting up optics and performing better tests of my software this week.

  1777   Wed Jul 22 11:18:49 2009 robConfigurationComputerssticky sliders

Quote:

Quote:

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

 Alberto, Rob,

we diagnosed the problem. It was related with sticky sliders. After a reboot of C1:IOO the actual output of the DAC does not correspond anymore to the values read on the sliders. In order to update the actual output it is necessary to do a change of the values of the sliders, i.e. jiggling a bit with them.

 I've updated the slider twiddle script to include the MC alignment biases.  We should run this script whenever we reboot all the hardware, and add any new sticky sliders you find to the end of the script.  It's at

 

/cvs/cds/caltech/scripts/Admin/slider_twiddle

  1776   Wed Jul 22 11:14:11 2009 AlbertoConfigurationComputerscomputers are down

Quote:

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

 Alberto, Rob,

we diagnosed the problem. It was related with sticky sliders. After a reboot of C1:IOO the actual output of the DAC does not correspond anymore to the values read on the sliders. In order to update the actual output it is necessary to do a change of the values of the sliders, i.e. jiggling a bit with them.

  1775   Wed Jul 22 11:08:36 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

I have built a version of the circuit with flying components; the completed circuit is shown in the attached picture. I built the circuit in segments and measured the transfer function after each segment to see whether it matched the LTSpice simulation after each step. The segments are shown in the circuit diagram.

After building the first segment, the measured transfer function looked pretty much the same as the simulated transfer function; it appears shifted in the attached plot, but this is because I didn't do a careful job of tuning at this point, and I'm relatively sure that I could have tuned it to match the simulation. After adding the second segment of the circuit, the measured and simulated transfer functions were similar in shape, but I was unable to increase the frequency of the peaks (through tuning) any more than what is shown in the plot (I could move the peaks so that their frequency was lower, but they are shown as high as they will go). When I added the final segment to complete the circuit, the measured and simulated transfer functions no longer had the same shape; two of the peaks were very close together and I was barely able to differentiate one from the other.

In order to understand what was happening, I tried making modifications to the LTSpice model to recreate the transfer function that was measured. I was able to create a transfer function that closely resembles the measured transfer function in both the circuit as of the 2nd segment and the completed circuit by adding extra inductance and capacitance as shown in red in the circuit diagram. The transfer functions simulated with these parasitic components are shown in red in both plots. While I was able to recreate the response of the circuit, the inductance and capacitance needed to do this were much larger than I would expect to occur naturally within the circuit (2.2uH, 12 pF). I'm not sure what's going on with this.

Attachment 1: BuiltCkt_Picture.png
BuiltCkt_Picture.png
Attachment 2: BuiltCkt2_Final.png
BuiltCkt2_Final.png
Attachment 3: 1stSegment.png
1stSegment.png
Attachment 4: 2ndSegment_ExtraL.png
2ndSegment_ExtraL.png
Attachment 5: Complete_ExtraL.png
Complete_ExtraL.png
  1774   Wed Jul 22 10:49:40 2009 AlbertoUpdatePSLMC Alignment Trend

Trying to track the MC positions back for a few days, it seems that the data hasn't been recorded properly for a while. Something happened yesterday after my boot fest and then the record got restored. Attached here are the readbacks showing the event for MC1.

Is anything wrong with the data record?

Attachment 1: 2009-07-21_MC-align-trend.pdf
2009-07-21_MC-align-trend.pdf
Attachment 2: 2009-07-22_mc-align-trend.pdf
2009-07-22_mc-align-trend.pdf
  1773   Wed Jul 22 09:04:10 2009 AlbertoConfigurationComputerscomputers are down

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

  1772   Wed Jul 22 01:57:19 2009 AlbertoConfigurationComputerscomputers are down

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

  1771   Tue Jul 21 18:46:47 2009 steveConfigurationComputerscomputers are down

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

  1770   Tue Jul 21 17:52:12 2009 JenneUpdateIOOMC_L flatlined

Quote:

[Clara, Jenne]

While Clara was working on her Wiener filtering and optimizing the locations of the accelerometers, she discovered that MC_L and MC_L_256 are totally flatlined.  I looked at them, and it looks like they've been dead since ~9:30pm-ish on Sunday night.  Bootfest-type activities shall commence shortly.

 Under Alberto's tutalage, I rebooted the whole vme set (iovme, sosvme, susvme1, susvme2), and after that MC_L was all good again.

  1769   Tue Jul 21 17:01:18 2009 peteDAQDAQtemp channel PEM-PETER_FE

I added a temporary channel, to input 9 on the PEM ADCU.    Beware the 30, 31, and 32 inputs.  I tried 32 and it only gave noise.

 

 

  1768   Tue Jul 21 15:32:47 2009 JenneUpdateIOOMC_L flatlined

[Clara, Jenne]

While Clara was working on her Wiener filtering and optimizing the locations of the accelerometers, she discovered that MC_L and MC_L_256 are totally flatlined.  I looked at them, and it looks like they've been dead since ~9:30pm-ish on Sunday night.  Bootfest-type activities shall commence shortly.

  1767   Tue Jul 21 13:55:08 2009 ClaraUpdatePEMGuralp Box Success!

There managed to be just enough 100 kOhm resistors to stuff all the "2" channels (VERT2, N/S2, E/W2) with the fancy low-noise resistors. The first six channels (VERT 1/2, NS 1/2, EW 1/2) are now completely done with the thin-film resistors, taking into account the changes that were made on the circuit diagram. I also replaced the C8 capacitor with the fancy Garrett ones and added capacitors on top of R4 and R13 (after painstakingly making sure that the capacitances are exactly the same for each pair) for the "2" channels. It looks like the capacitors on the "1" channels are the cheaper ones. I will compare the noise measurements later to see if there is any difference - if so, I can replace those as well (although, we're out of the 1 uF capacitors needed for C8).

Speaking of, we are now out of or very low on several types of the Garrett resistors/capacitors: 1 uF, 1kOhm, 100 Ohm, 14.0 Ohm, and 100 kOhm. I left the specifics on Steve's desk so that more can be ordered for the eventual time when the third set of channels needs to be restuffed.

  1766   Tue Jul 21 01:11:30 2009 DmassAoGComputersAlarms going off

I came into the 40m to sign things out briefly then swiftly return them, and the alarms were going off on op540m at 1am.

 

The cat and donkey? were making much noise.

  1765   Mon Jul 20 17:06:29 2009 JenneUpdatePEMGuralp Box Fail

Quote:
That's terrible: R5 & R8 should definitely be 100 Ohm and not 100kOhm. 100k would make it a noise disaster. They should also be metal film (from the expensive box, not from the standard box). This is the same for all channels so might as well stuff them.

The circuit diagram between TP3 and TP4 appears to be designed to make the whitening not work. That's why R6 & R7 should be 100k. And R2 should be metal film too.

Basically, every time we want good low frequency performance we have to use the metal film or metal foil or wirewound resistors. Everything else produces a lot of crackling noise under the influence of DC current.

I'm also attaching the voltage and current noise spectra for the AD620 from the datasheet. These should allow us to compare our measurements to a reasonable baseline.


While we're comparing things to other things, Ben Abbott just emailed me his measurement of the AD620 from back in the day. Clara's going to use this along with the specs to make sure that (a) we're not taking crazy measurements and (b) our AD620s aren't broken and in need of replacement. In this plot, we're looking at the GOLD trace, which has the AD620 set up with a gain of 10, which is how our AD620's are set up in the Guralp breakout box.

Just picking a single point to compare, it looks like at 1Hz, Ben saw ~130dBVrms/rtHz. Converting this to regular units [ 10^(#dB/20)*1Vrms = Vrms ], this is about 3*10^-7 Vrms. That means that Clara's measurements of our AD620 noise is within a factor of 2 of Ben's. Maybe the way we're connecting them up just isn't allowing us to achieve the ~50nV/rtHz that is claimed.
Attachment 1: AD620noise_BenAbbott.pdf
AD620noise_BenAbbott.pdf
  1764   Mon Jul 20 12:35:21 2009 robConfigurationLockingalignment biases funny

I found the alignment biases for the PRM and the SRM in a funny state.  It seemed like they had been "saved" in badly misaligned position, so the restore scripts on the IFO configure screen were not working.  I've manually put them into a better alignment.

  1763   Mon Jul 20 10:35:06 2009 steveUpdateVACUPS batteries replaced

 APC Smart-UPS (uninterruptible power supply) batteries RBC12 replaced at 1Y8 vacuum rack.

Their life span were 22 months.

  1762   Sun Jul 19 22:38:24 2009 robOmnistructureGeneralWeb screenshots aren't being updated

Quote:

Before heading back to the 40m to check on the computer situation, I thought I'd check the web screenshots page that Kakeru worked on, and it looks like none of the screens have been updated since June 1st.  I don't know what the story is on that one, or how to fix it, but it'd be handy if it were fixed.

 Apparently I broke this when I added op540m to the webstatus page.  It's fixed now.

  1761   Sat Jul 18 19:49:48 2009 ranaUpdatePEMGuralp Box Fail
That's terrible: R5 & R8 should definitely be 100 Ohm and not 100kOhm. 100k would make it a noise disaster. They should also be metal film (from the expensive box, not from the standard box). This is the same for all channels so might as well stuff them.

The circuit diagram between TP3 and TP4 appears to be designed to make the whitening not work. That's why R6 & R7 should be 100k. And R2 should be metal film too.

Basically, every time we want good low frequency performance we have to use the metal film or metal foil or wirewound resistors. Everything else produces a lot of crackling noise under the influence of DC current.

I'm also attaching the voltage and current noise spectra for the AD620 from the datasheet. These should allow us to compare our measurements to a reasonable baseline.
Attachment 1: Picture_1.png
Picture_1.png
Attachment 2: Picture_2.png
Picture_2.png
Attachment 3: AD620.pdf
AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf
  1760   Fri Jul 17 18:04:54 2009 ClaraUpdatePEMGuralp Box Fail

I've been trying for most of the week to get noise measurements on the output of the Guralp box as well as scross the AD640 chip. The measurements haven't really been making sense, and, being at a loss as to what else I should try, I decided to redo the resistors on the N/S 2 and E/W 2 channels. (I had been comparing the VERT1 and VERT2 channels, as VERT1 has been restuffed and VERT2 has not.) I don't need all three of the second set of channels to do more measurements, so it seemed like a good use of time.

The first thing I noticed was that the VERT2 channel was missing two resistors (R24 and R25). I probably should have noticed this sooner, as they are right by the output points I had been measuring across, but it didn't occur to me that anyone did anything to the VERT2 channel at all. So, probably the measurements on VERT2 are no good.

VERT2_missing_resistors.png

Note the existence of 100 kOhm resistors on the top channel, and none on the bottom channel (VERT2).

 

Then, while I was soldering in some 100 Ohm resistors, I happened to notice that the resistors I was using had a different number (1001) on them than the corresponding ones on the already redone channels (1003). I checked the resistance, and the ones on the already redone channels turned out to be 100 kOhm resistors, rather than 100 Ohms. So, I double checked the circuit diagram to make sure that I had read it correctly, and there were a number of resistors that had been relabeled as 100 Ohms and several relabeled as 100 kOhms. On the board, however, they were ALL 100 kOhms. Clearly, one of them is wrong, and I suspect that it is the circuitboard, but I don't know for sure.

resistors_diagram.png

resistors_board.png

The diagram clearly shows that R6 should be a 100k resistor, while R5 and R8 should be 100 Ohm resistors, but they are all the same (100k) on the board. I suspect this may have something to do with larger-than-expected noise measurements. But, it's possible the diagram is wrong, not the board. In any case, I didn't really know what to do, since I wasn't sure which was right, so I just replaced all the resistors I was sure about and removed the 100k and 100 Ohm resistors without replacing them with anything. Incidentally, the box of 100kOhm resistors seems to be missing, so I wouldn't have been able to finish those anyway.

  1759   Thu Jul 16 14:54:05 2009 robUpdateLockingMC_F channel dead

Quote:

Quote:

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

 Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack).  Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.

 This has been fixed by one of the two most powerful & useful IFO debugging techniques: rebooting.  I keyed the crate in 1Y2.

  1758   Thu Jul 16 14:41:38 2009 robUpdateLockingMC_F channel dead

Quote:

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

 Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack).  Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.

Attachment 1: IOO_ICS_0_15.png
IOO_ICS_0_15.png
Attachment 2: IOO_ICS_15_32.png
IOO_ICS_15_32.png
  1757   Thu Jul 16 10:52:58 2009 ClaraUpdatePEMSingle Channel TRS-RNC Cable

I made and tested a female-to-female TRS(audio)-RNC cable. It only has a single channel, so it won't work for stereo speakers or anything, but I should only need one speaker for testing the microphones. The tip of the plug is the signal, the sleeve is ground, and the ring is null.

  1756   Thu Jul 16 09:49:52 2009 AlanUpdateComputersfb40m

Quote:

The fb40m just went out of order with status indicator number 8

It recovered on its own five minutes later.

 Backup script restarted, backup of trend frames and /cvs/cds is up-to-date.

 

  1755   Thu Jul 16 01:00:56 2009 AlbertoUpdateLockingPD9 aligned

Quote:

Quote:

Since lately the alignment of the input beam to the interferometer has changed, I went checking the alignment of the beam on the photodiodea. They were all fine except for pd9, that is AS DD 199. Here the DC is totally null. The beam seems to go right on the diode but the scope on the PD's DC output shows no power. This is really strange and bad.

After inspecting PD9 with the viewer and the cards, the beam looks like it is aligned to the photodiode althought there is no signal at the DC output of the photodetector. So I checked the spectrum for PD9_i and Q (see attachments) and it seems that those channels are actually seeing the beam. I'm going to check the alignemtn again and see the efefct on the spectra to make sure that the beam is really hitting the PD.

 

 I aligned PD9. Here are the spectra confirming that.

p.s.
Ants, theyre everywhere, even inside the AS table. They're taking over the lab, save yourself!
Attachment 1: 2009-07-15_PD9spectrumPDF02.pdf
2009-07-15_PD9spectrumPDF02.pdf
  1754   Wed Jul 15 18:35:11 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

Using FET probes, I was able to measure a transfer function that looks a little more like what I expected. There are only two peaks, but I think this can be explained by a short between the two capacitors (and two tunable capacitors) in the LC pairs, as shown (in red) in the circuit diagram attached. The measured transfer function (black), along with the simulated transfer functions with (red) and without (blue) the short are shown in the attached plot. The measured transfer function doesn't look exactly like the simulated transfer function with the short, but I think the difference can be explained by stray impedances.

Attachment 1: BuiltCkt1_Final.png
BuiltCkt1_Final.png
Attachment 2: BuiltCkt1_TransferFunctions.png
BuiltCkt1_TransferFunctions.png
  1753   Wed Jul 15 18:22:15 2009 KojiUpdateCamerasRe: GigE Phase Camera

Quote:

Koji recommended that we use the optical setup pictured below.  Although it uses fewer optics, I can't think of a way to test the phase camera using this configuration because any modulation of the wavefront with a lens or whatever would be automatically corrected for in the PLL so I think I'll have to stick with the old configuration.

I talked with Zach. So this is just a note for the others.

The setup I suggested was totally equivalent with the setup proposed in the entry http://131.215.115.52:8080/40m/1721, except that the PLL PD sees not only 29.501MHz, but also 1kHz and 59.001MHz. These additional beating are excluded by the PD and the PLL servo. In any case the beating at 1kHz is present at the camera. So if you play with the beamsplitter alignment you will see not only the perfect Gaussian picture, but also distorted picture which is resulted by mismatching of the two wave fronts. That's the fun part!

The point is that you can get an equivalent type of the test with fewer optics and fewer efforts. Particularly, I guess the setup would not be the final goal. So, these features would be nice for you.

  1752   Wed Jul 15 17:18:24 2009 JenneDAQComputersDAQAWG gone, now back

Yet again, the DAQAWG flipped out for an unknowable reason.  In order of restart activities listed on the Wiki, I keyed the crate and nothing really happened, then I hit the physical reset button and nothing happened, and then I did the 'telnet....vmeBusReset', and a couple minutes later, it was all good again.

  1751   Wed Jul 15 14:42:31 2009 ZachUpdateCamerasGigE Phase Camera

Lately, I have been able to externally trigger the camera using a signal generator passing through the op-amp circuit that I built.  The op-amp circuit stabilizes the jitter in the sine wave from the signal generator and rectifies the wave.  I wrote the calculations into the code allowing me to find the phase and amplitude from the images I take.  I still need to develop code that will plot these arrays of phase and amplitude.

The mysterious dark band at the top of the ccd images continues to defy explanation.  However, I have found that it only appears for short exposure times even when the lens is completly covered.  During the next couple of days, I will try to write a routine to correct for this structure in the dark field.

Koji recommended that we use the optical setup pictured below.  This configuration would require fewer optics and we would have to rely on slight misalignments between the carrier and reference beams to test the effectiveness of the phase camera instead of a wavefront-deforming lens.

Attachment 1: fig1koji.pdf
fig1koji.pdf
  1750   Wed Jul 15 12:44:28 2009 Chris ZimmermanUpdateGeneralWeek 4/5 Update

I've spent most of the last week working on finishing up the UCSD calculations, comparing it to the EUCLID design, and thinking about getting started with a prototype and modelling in MATLAB.  Attached is something on EUCLID/UCSD sensors.

Attachment 1: Comparison.pdf
Comparison.pdf Comparison.pdf
  1749   Wed Jul 15 12:14:08 2009 AlbertoUpdateLockingphotodiode alignment check

Quote:

Since lately the alignment of the input beam to the interferometer has changed, I went checking the alignment of the beam on the photodiodea. They were all fine except for pd9, that is AS DD 199. Here the DC is totally null. The beam seems to go right on the diode but the scope on the PD's DC output shows no power. This is really strange and bad.

After inspecting PD9 with the viewer and the cards, the beam looks like it is aligned to the photodiode althought there is no signal at the DC output of the photodetector. So I checked the spectrum for PD9_i and Q (see attachments) and it seems that those channels are actually seeing the beam. I'm going to check the alignemtn again and see the efefct on the spectra to make sure that the beam is really hitting the PD.

 

Attachment 1: 2009-07-15_PD9spectrumPDF.pdf
2009-07-15_PD9spectrumPDF.pdf
  1748   Wed Jul 15 12:11:17 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

This week I've been working on testing the first version of the prototype circuit. Initially, I tested the circuit that I built last week, which had resistors in the place of the transformer. The magnitude and phase of the transfer function, as measured by the Agilent 4395A, are shown in the attached plot (first plot, MeasuredTransferFunction_R.jpg). The transfer function doesn't look like the simulated transfer function (second plot, BuiltCkt_ExpectedResponse.png), but I think I see the three peaks at least (although they're at the wrong frequencies). I spent some time trying to recreate the actual transfer function using LTSpice, and I think it's reasonable that the unexpected response could be created by extra inductance, resistance, capacitance and interaction between components.

When the transformer arrived  yesterday, I replaced the resistors in the circuit with the transformer, and I have measured the following response (last plot, MeasuredTransferFunction.jpg). The gain is much lower than for the circuit with the resistors; however, I am still trying to track down loose connections, since the measured transfer function seems very sensitive to jiggled wires and connections.

Meanwhile, the parts for a flying-component prototype circuit have been ordered, and when they arrive, I'll build that to see if it works a little better.

Attachment 1: MeasuredTransferFunction_R.jpg
MeasuredTransferFunction_R.jpg
Attachment 2: BuiltCkt_ExpectedResponse.png
BuiltCkt_ExpectedResponse.png
Attachment 3: MeasuredTransferFunction.jpg
MeasuredTransferFunction.jpg
  1747   Wed Jul 15 11:38:31 2009 robUpdateLockingMC_F channel dead

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

Attachment 1: mcfdead.png
mcfdead.png
  1746   Wed Jul 15 08:59:30 2009 steveUpdateComputersfb40m

The fb40m just went out of order with status indicator number 8

It recovered on its own five minutes later.

  1745   Tue Jul 14 17:48:20 2009 JenneOmnistructureEnvironmentRemoval of the cold air deflection device for the MOPA chiller

Quote:

Quote:

Quote:
Around 2 PM today, I removed the blue flap which has been deflecting the cold air from the AC down into the laser chiller.
Let's watch the laser trends for a few days to see if there's any effect.


Alberto has moved us to stage 2 of this experiment: turning off the AC.

The situation at the control room computers with the AC on minus the blue flap is untenable--it's too cold and the air flow has an unpleasant eye-drying effect.


I turned the AC back on because the temperature of the room was going up so also that of the laser chiller.


I reinstalled the blue-flap technology on the AC, because the MOPA power was dropping like a rock. A light-ish rock since it wasn't going down too fast, but the alarms started going a little while ago because PMC trans was too low, because the power was getting a little low. The laser water chiller is reading 21.97C, which is higher than it normally does/did before the AC shenanigans (It usually reads 20.00C).

Attached is a look-back of 18 hours, during which you can see in the AMPMON the time that Rana removed the blue flap around 2pm yesterday and the AMPMON changes a little bit, but not drastically, the time around 11pm when the AC was turned off, and AMPMON goes down pretty fast, and about 12:30am, when Alberto turned the AC back on, and AMPMON starts to recover. I think that the AMPMON starts to go down again in the morning because it's been crazy hot here in Pasadena, so the room might be getting warmer, especially with the laser chiller-chiller not actively chilling the laser chiller (by not being pointed at the water chiller), so the water isn't getting as cold, and the HTEMP started to go up.

In the last few minutes of having put the blue flap back on the AC, the laser chiller is already reading a lower temperature, and the AMPMON is starting to recover.
Attachment 1: ACeffectonLASER2.png
ACeffectonLASER2.png
  1744   Tue Jul 14 16:31:46 2009 ClaraUpdatePEMFrequency Response of Microphone (Bonnie)

So, I actually took these measurements last week, but I didn't get around to making nice plots and things until now. I figured the time while I wait for the spectrum analyzer to do its thing was a good time.

Having been unable to locate the SR785 and also unsure how to connect it to a computer speaker (and also unable to find a free one), I downloaded a demo of a function generator onto my computer and just used that. (Same thing I used to do the swept sine that created the frequency power response plots I posted last week.) I set the program to a number of different frequencies and had the other end of the cable hooked into the oscilloscope to see a) if I could pick out the frequency and b) see how the magnitude of the microphone output varied with the frequency.

The first set of measurements I took, I didn't realize that I could increase the output power of the function generator. Because the generated sound at the default setting was relatively quiet, the oscilloscope traces were pretty chaotic, so I usually froze the trace so that I could look at it better. I ended up with a lot of weird jumps in the magnitude, but I later realized that there was a lot of beating going on at some frequencies, and the amplitude changes were probably much more drastic for the -20 dB sounds than the 6 dB sounds, since it was closer in amplitude to the surrounding noises. So, I've included that data set in my plots for the sake of completeness, but I'm pretty sure that it is useless.

Once I realized I could increase the power output for the signal generator, I took a set of data with and without the voltage divider at 6 dB. There was a cluster of frequencies that showed significant beating around 1700-3000 Hz in the data WITH the voltage divider, but I did not see any clear beating in the data WITHOUT. In the plots, I simply plotted up the highest and lowest amplitudes I measured for the frequencies with significant beating, since it was obviously hard to tell what the amplitude would have been without any background noise. In the w/o volt. div. set, although I didn't see any obvious beat patterns, the measured amplitudes did jump slightly at the frequencies that showed beats with the voltage divider. So, perhaps I was just not seeing them, but they influenced my amplitude measurements? I'm not sure if it would be possible for the voltage divider itself to cause beat frequencies.

 (Note: the amplitudes measured were from zero to peak, as the oscilloscope I was using wouldn't show a big enough vertical range to easily measure the peak-to-peak voltage difference.)

I've attached two plots of my measurements. One has a regular x-scale and includes all the measurements. The second has a logarithmic x-scale and omits the 20 Hz points. I had some troubles being able to pick out the 20 Hz signal on the oscilloscope... I don't know if my computer speakers just don't work well at that frequency or what, but either way, those points seemed highly suspect, and omitting them from the log plot allowed me to spread things out more.

One thing I'm not sure about is the 3000 Hz point. It was one of the ones with a beat frequency (~130 Hz), and the amplitudes were pretty low. The corresponding point from the non-voltage-divider data set is also low. So, I'm not sure what's happening there.

The one thing that I do think is quite clear is that the 1000 Hz drop-off in power when the microphone is connected to the ADC has nothing to do with the voltage divider. Beat issues aside, the shapes are very similar (pay no attention to the absolute scale... obviously, the voltage responses with and without the voltage divider were very different, and I just scaled them to fit in the same plot).

Update: Jenne pointed out that I was not absolutely clear about the voltage scale in my plots. The GREEN and BLUE points are on a mV scale, and the RED points are on a 10mV scale. I should probably redo the plots in Matlab in eventuality, since Excel is hard to use if you want to do anything that is not extremely basic with your plots, but this was my solution for the time being. So, the fact that the RED points, which are the data taken WITHOUT the voltage divider, are lower than the GREEN ones does not in any way indicate that I measured lower voltages when the voltage divider was not used.

Also, a to do list:

- Many of the beat frequencies I picked out were veeeeery slow, indicating that something is going at a frequency that is very close to the arbitrary frequencies I chose to sample, which is a little strange. That, combined with the fact that I saw clear beats with the voltage divider but not without leads me to believe that it may be worth investigating the frequency response of the voltage divider itself.

- Redo the measurements near the anomalous 3000 Hz point with a higher density of sampled frequencies to try to see what the heck is going on there.

Attachment 1: Bonnie_fres_regplot.pdf
Bonnie_fres_regplot.pdf
Attachment 2: Bonnie_fres_logplot.pdf
Bonnie_fres_logplot.pdf
  1743   Tue Jul 14 14:54:19 2009 steveConfigurationComputersfb40m2 in 1Y6

Alex and Steve,

SunFire x4600 ( not  MEGATRON 2 , it is fb40m2 ) and JetStor ( 16 x 1 TB drives ) were installed on side rails at the bottom of 1Y6

We cleaned up the fibres and cabling in 1Y7 also

ELOG V3.1.3-